OpenStudy (anonymous):

Find g º f if f(x) = 2x - 1 and g(x) = x + 2. 2x - 3 2x + 1 2x - 1 2x + 3

OpenStudy (anonymous):

My answer is b

OpenStudy (muzzack):

g and f you mean

OpenStudy (muzzack):

yes i think its b

OpenStudy (anonymous):

That is what the question said, but I think so

OpenStudy (anonymous):

(g∘f)(x) = g(f(x)) = f(x) + 2 = (2x - 1) + 2 = 2x + 1

OpenStudy (anonymous):

Thank you both!